OpenStudy (anonymous):

Solve triangle ABC given that A=56degrees, B=59degrees, b=68 a.) C=245degrees, a =65.77, c=71.9 b.) C=65degrees, a=70.31, c=74.34 c.) C=245degrees, a = 70.31, c=74.34 d.)C= 65degrees , a=65.77, c=71.9

OpenStudy (johnweldon1993):

Have you used the law of sines before?

OpenStudy (anonymous):

I haven't, no :/

OpenStudy (johnweldon1993):

Well it states that \[\frac{ Sin A }{ a } = \frac{ SinB }{ b } = \frac{ SinC }{ c }\] Where A,B, and C are angles...and a, b, and c are side lengths....you can use this to solve for the side of A that you need.... to use it...you need at least 1 side length...and its corresponding angle...which you have...

OpenStudy (johnweldon1993):

Plug in your angle 'B' and side length 'b' and angle 'A' into this and solve for 'a' \[\frac{ Sin(56) }{ a } = \frac{ Sin(59) }{ 68 }\]

OpenStudy (johnweldon1993):

Does that make sense?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Thank you @johnweldon1993